OpenStudy (anonymous):

I will medal!!! Which of the following ordered pairs could be placed in the table and still have the relation qualify as a linear function? Explain how you got it! Input (x) Output (y) 1 12 3 9 5 6 ? ? (7, 3) (7, 10) (3, 3) (3, 10)

OpenStudy (godlovesme):

u see that the x is increasing by ??/ and the y is increasing by ??? each time @Plm9

OpenStudy (godlovesme):

@Plm9 please participate :)

OpenStudy (anonymous):

x is by 2 but y is by 3 decreasing

OpenStudy (godlovesme):

so add 2 to 5 and subtract 3 from 6

OpenStudy (godlovesme):

@Plm9 still there?

OpenStudy (anonymous):

yes can you explain to me how you got that?

OpenStudy (godlovesme):

what's 5+2 = ? 6-3 = ?

OpenStudy (anonymous):

5/2= 2.5 6-3=3

OpenStudy (godlovesme):

i said 5+2 not 5/2 :P

OpenStudy (anonymous):

7

OpenStudy (godlovesme):

so, x will be 7 and y will be 3 (x,y) => (7,3) got it ? @Plm9

OpenStudy (anonymous):

thank you very much Godlovesme that was very help full!

OpenStudy (godlovesme):

u r welcome :) <3